Handbook of Coding Theory

V. S. Pless, W. C. Huffman, R. A. Brualdi, editors

1 Algebraic Coding

1. An Introduction to Algebraic Codes – V. S. Pless, W. C. Huffman, and R. A. Brualdi 2. Self-dual Codes – J. H. Conway and N. J. A. Sloane

3. Combinatorial Constructions – V. S. Pless

4. Covering Radius – R. A. Brualdi, S. Litsyn, and V. S. Pless

5. Codes from Algebraic Geometry –T. Høholdt, J. H. van Lint, and R. Pellikaan

6. Asymptotic Bounds – V. I. Levenshtein

7. Optimal Codes and Tables – A. Brouwer and N. J. A. Sloane

8. Algebraic Theory of Convolutional Codes – R. J. McEliece

9. Open Problems in Cyclic Codes – P. Charpin

2 Codes and …

1. Groups – W. C. Huffman

2. Number Theory – I. Honkala and A. Tietäväinen

3. Designs – V. D. Tonchev

4. Finite Geometries – E. F. Assmus Jr. and J. D. Key

5. Quadratic Forms (including Representations) – H. N. Ward

6. Association Schemes – P. Camion

7. Lattices – J. H. Conway and N. J. A. Sloane

8. Crytography – H. van Tilborg

9. Sequences and Arrays – T. Helleseth and V. J. Kumar

3 Applications

1. Decoding Algorithms – R. E. Blahut

2. Concatenated Codes – I. Dumer

3. Trellis Codes – A. Vardy

4. Codes for Magnetic Recording – B. H. Marcus, R. Roth, and P. H. Siegel

5. Deep Space Applications of Codes – S. Wicker

6. Array Codes – M. Blaum, P. Farrell, and H. Van Tilborg